1. Coronado Company purchases equipment on January 1, Year 1, at a cost of $540,000. The asset is expected to have a service life of 12 years and a salvage value of $48,600.

Compute the amount of depreciation for each of Years 1 through 3 using the straight-line depreciation method. (Round answers to 0 decimal places, e.g. 5,125.)

Depreciation for Year 1

$enter a dollar amount rounded to 0 decimal places

Depreciation for Year 2

$enter a dollar amount rounded to 0 decimal places

Depreciation for Year 3

$enter a dollar amount rounded to 0 decimal places

Compute the amount of depreciation for each of Years 1 through 3 using the sum-of-the-years'-digits method.

Depreciation for Year 1

$enter a dollar amount

Depreciation for Year 2

$enter a dollar amount

Depreciation for Year 3

$enter a dollar amount

2. Windsor Corporation purchased a new machine for its assembly process on August 1, 2020. The cost of this machine was $136,500. The company estimated that the machine would have a salvage value of $17,700 at the end of its service life. Its life is estimated at 5 years, and its working hours are estimated at 20,000 hours. Year-end is December 31.

Compute the depreciation expense under the following methods. Each of the following should be considered unrelated.

(a)

Straight-line depreciation for 2020

$enter a dollar amount

(b)

Activity method for 2020, assuming that machine usage was 700 hours

$enter a dollar amount

(c)

Sum-of-the-years'-digits for 2021

$enter a dollar amount

(d)

Double-declining-balance for 2021

$enter a dollar amount

3. Kingbird Corporation acquires a coal mine at a cost of $484,000. Intangible development costs total $121,000. After extraction has occurred, Kingbird must restore the property (estimated fair value of the obligation is $96,800), after which it can be sold for $193,600. Kingbird estimates that 4,840 tons of coal can be extracted.

If 847 tons are extracted the first year, prepare the journal entry to record depletion.

Account Titles and Explanation

Debit

Credit

enter an account title

enter a debit amount

enter a credit amount

enter an account title

enter a debit amount

enter a credit amount
Answer #1

1.
Straight line Depreciation = (Original Value - Salvage Value) / Useful Life
= ($540000-48600)/12 = $40950 per year

Depreciation Year 1 $          40,950
Depreciation Year 2 $          40,950
Depreciation Year 3 $          40,950

Sum of Years Digit Method
Sum of Years = 1+2+3+4+5+6+7+8+9+10+11+12 = 78

Depreciation = (Original Value - Salvage Value) / Sum of Years x Remaining Life

Depreciation Year 1 $          75,600 =(540000-48600)/78*12
Depreciation Year 2 $          69,300 =(540000-48600)/78*11
Depreciation Year 3 $          63,000 =(540000-48600)/78*10

Note : As per HOMEWORKLIB RULES we are supposed to do one question completely, so kindly post other questions separately.